# Are We Asking the Right Questions?: Designing for Community Stakeholders’ Interactions with AI in Policing

### 08 Feb 2024 v1

@@ 4.2. Participants’ Interactions With Algorithmic Crime Mapping (RQ2)
We conducted in-lab activities with our participants where we first gave them a brief overview of the crime mapping application and explained to them how they could change parameters to produce different maps. Next, we asked the participants to estimate the number of hotspots on the maps based on their knowledge of the city.
@@
We learned that LEAs used their hyperlocal knowledge of the different intersections (within neighborhoods) in the city to identify epicenters for criminal activity, however, in doing so, significantly overestimated the number of hotspots. For instance, an LEA who was interacting with the crime map shared that he was looking at all the major intersections in the vicinity of hotspots and tying criminal activity to his street-level knowledge of those intersections -
@@
LEAs could have been overestimating hotspots for several reasons. Evidence in criminology literature indicates that police officers may overestimate the prevalence of crime in certain neighborhoods or communities (buil2021accuracy, ). Here, the perception of crime can be influenced by several factors such as the nature of policing work, media coverage of crime, and perceptions about low-income and minority communities (smith2007explaining, ). There is also evidence that indicates that micro-level crime analysis as undertaken by LEAs (i.e., assessing specific intersections) is prone to a larger risk of selection and reporting bias as compared to crimes aggregated and assessed at a larger scale. However, in our conversations with the LEAs, we also learned that they did not draw clear distinctions between violent crimes and quality-of-life crimes. They often referred to the fact that different types of crimes can be interlinked and at times quality-of-life crimes can escalate to more serious crimes. That is, they drew on their domain expertise to focus on several hyperlocal hotspots instead of a general hotspot covering a neighborhood that a squad car could easily patrol. This finding was further supported by the next activity where we asked participants to place green circles on the map that demarcate the area that one patrol car should cover.
@@
During our observations of participants’ interactions with different crime maps, we recorded whether participants changed the parameters to create different maps and their responses to these different maps. Please see Appendix B for results of the quantitative analysis. Both technical participants and community members interacted with parameters and generated different maps. Some participants gamified the process for themselves and started to guess how changing certain parameters would affect the map. Technical participants shared that they had seen static geographical heatmaps before and knew how to generate them by writing code but had never interacted with an interactive map. T9 shared -
@@
Several community members appreciated the opportunity to interact with such an application and be able to understand how resources might be allocated using such tools. They shared that such transparency, even though carried out as a lab activity, helped them develop trust. For instance, a community member shared -
@@
Several community members also shared that they liked the easier maps because they provided them with bigger circles (i.e., hotspots) that gave them a quick overview of the neighborhoods and would be a useful communication and awareness tool for the community.
The results, regarding LEAs, were quite surprising for the research team. We had anticipated that LEAs would be most likely to interact with the parameters, generate different maps, and connect the findings to street-level occurrences over different periods. However, none of the LEAs changed parameters and spent a significant amount of time interpreting the first map presented to them. Instead of using the crime maps to augment their knowledge of geospatial and temporal characteristics of different crimes, LEAs quickly started connecting their street-level domain knowledge to the hotspots on the map, i.e. - the first piece of information presented to them. Instead of using the tool as a decision aid that would augment their knowledge, they used it as a means to validate what they already knew about certain neighborhoods, sections of highways, and intersections. When urged to change parameters and create a different map, an LEA who had experience working with crime maps stated -
@@
The LEA makes an interesting point here regarding the data that we further delve into in the next subsection. However, for the most part, we witnessed that LEAs exhibited a strong anchoring bias where the default map (i.e., the ‘anchor’) provided a ‘satisfactory outcome’ that seemed to align with what they already knew about the city. Anchoring bias is a cognitive bias that refers to the tendency of practitioners to rely too heavily on the first piece of information encountered when making decisions and has been observed across several domains (furnham2011literature, ; azzopardi2021cognitive, ; saposnik2016cognitive, ). However, as we discuss in the next section (see Section 4.3), the LEAs asked for features (e.g., the ability to draw districts on the map, see spatiotemporal relationships between different types of crime, and use information from on-the-ground investigations) that would allow them to derive more context-specific information.
@@
We wanted to assess participants’ level of engagement while interacting with the application and their mental workload as a result of this interaction. To accomplish this, we conducted the NASA-TLX survey and assessed participants on four factors (i.e., mental demand, temporal demand, performance, and frustration level). To assess the differences in scores across participants and their significance, we conducted statistical tests the details of which are provided in Appendix C.
@@
The third parameter we considered was performance (PF) which determines the success of the users in carrying out their duties and how satisfied they were with the results of their work. Both the LEAs and technical participants scored high on this parameter which depicts that these participants were confident about their results. On the other hand, community members’ scores were in the medium range implying that they were uncertain about their results. As depicted in the previous section both the LEAs and technical participants had some experience working with crime maps or geospatial heatmaps. However, technical participants interpreted maps based on their understanding of heatmaps whereas LEAs interpreted maps based on their street-level domain knowledge.
@@
Taken together, the results for LEAs are surprising. The NASA-TLX results show that LEAs were deeply invested in interpreting the maps presented to them (i.e., they scored high on mental and temporal demand), considered that they performed well when interpreting the hotspots and placing squad cars (i.e., scored high on performance) and did not find the task to be frustrating (i.e., scored low on frustration). However, as noted previously, none of the LEAs changed parameters to generate different maps. Here, LEAs were more interested in posthoc analysis and interpreting the information provided to them instead of generating new information that may help them uncover new underlying factors that they had not considered before in their practice.
4.3. Participants’ Needs Regarding the Ethical Design and Use of Algorithmic Crime Mapping (RQ3)
@@
Several community members and technical participants found utility in algorithmic crime mapping as a public resource, however, questioned the underlying motivation of police departments themselves. Much of our conversations in this regard centered around ongoing policing practices where they questioned the trustworthiness of police reports by drawing on cases of police brutality where the police reports did not match the video evidence shared by bystanders on social media. For instance, participant C21 shared -
”From my personal experience, it seems that in black neighborhoods police are stopping more drivers per month per hour for all kinds of things, and for things that they don’t stop white drivers in the suburbs for. There are penalties for more and more people for using illegal drugs in the city, but people living in the suburbs use the same drug. But they get off with a warning and they think they are somehow different and too good for prosecution and they get some kind of a better deal. So there is a wide widespread perception in society, that law enforcement is not really fair. And on one hand, I would like to believe ’no, it really is fair’. But the evidence is coming up more and more in recent years. Like with the development of the cell phone, people take photographs and take videos and it doesn’t match the police reports. Before we had this technology, and some person who was arrested would say that the police beat him up and we would hear the police officer say that ’he was resisting arrest’.” - C21, Community Member
This concern was shared by several community members who believed that such technologies could lead to racial profiling and amplify racial disparities. Participants shared examples of cases where the defendants were wrongfully convicted by association and forced into plea deals. Here, technical participants also brought into question the objectivity of the data analysis process itself and the need for public transparency and accountability mechanisms that keep such technologies in check. For instance, participant T5 shared the following when discussing the role of a data analyst working in a crime analysis lab -
@@
This concern was also echoed by other technical participants who shared examples of how data can be tweaked and cherry-picked and drew attention to the subjective nature of data science practice, an ongoing concern that has been raised in prior work (mothilal2024towards, ; das2024colonial, ; passi2017data, ; passi2020making, ). Finally, as discussed in Section 4.1, participants shared examples of differences in policing practices in urban and suburban neighborhoods, examples of speed traps, and people in low-income neighborhoods being set up for petty crimes to draw attention to policing practices that unethically target vulnerable communities instead of protecting and serving them. Here, community members sought accountability from police departments regarding street-level policing practices, whereas, technical participants wanted more transparency and accountability regarding data science practices that drive algorithmic crime mapping and data-driven policing.
@@ 4.3.2. Tensions Around the Ethical Basis of Data Collection
Several community members discussed the need for trustworthiness and accountability in regard to data collection practices and discussed instances of falsehoods in police reports, differences in data that is available about urban versus suburban communities, and policing practices designed to target people (e.g., speed traps). Here, some participants also recognized that several government agencies collect data about citizens while conducting their daily operations but wanted to ensure that it is collected lawfully and with some oversight from the judicial system. For instance, a participant shared -
@@
This concern was also echoed by LEAs who shared that it was necessary to collect data legally and ethically otherwise it could not be used for investigations and would not be admissible in court during legal proceedings. An LEA shared -
@@
Here, the LEA provides an example of the police department getting access to people’s cell phone data in a specific neighborhood, however, states that they must make their case before a judge and get a court order approved before receiving access. LEAs also shared that their real-time crime mapping technology was based on service calls and not police reports and that people’s biases played a significant role regarding the calls they receive and investigate -
@@
Several LEAs recognized the concerns surrounding racial profiling but drew attention to the nature of their work where they were required to investigate service calls that were assigned to them as well as patrol the zones they were assigned to. Similarly, several LEAs shared the nuances of their street-level labor that we discuss next.
@@ 4.3.3. Nuances of Policing Practices
”I never knew the algorithm that was used - that was never shared. There were a couple of occasions where me and my team would be given a packet of information about a specific crime that was occurring in our area of responsibility, and sometimes the predictive information that was given to us wasn’t consistent with what we knew as law enforcement officers just being in the neighborhood.” -L5, Law Enforcement Agent
@@
Here, it is interesting to note that both the technical participants and LEAs drew attention to the nature of data science practice. Technical participants discussed the elastic and subjective choices that data scientists make during analysis, whereas, LEAs discussed the need for data scientists to understand the nature of street-level discretionary work that they undertake. These themes have been observed across different contexts where domain experts have drawn attention to nuances of their professional practice that cannot be accounted for by quantitative data, and are generally unobservable to people outside the profession but significantly impact the decision-making process (kawakami2022improving, ; watkins_art, ; spektor2023designing, ).
@@ 5. Discussion
5.1. Eliciting Feedback from Community Members to ML Problem Formulation
@@
Early deliberations regarding the development and use of AI systems often omit impacted stakeholders because of their lack of technical or data science knowledge which supposedly limits them from providing meaningful feedback that can inform systems design (kuo2023understanding, ; delgado2023participatory, ). However, as our results highlight, community members, irrespective of technical knowledge, are able to provide detailed feedback regarding whether the technical formulation of a problem is even socially relevant, draw attention to systemic and structural issues that might be invisible to AI developers, and help uncover the downstream impact of such systems on their communities. That is, impacted stakeholders have a critical role to play at the earliest stages of machine learning problem formulation where data scientists translate broad and nondescript objectives into tractable problems that machine learning can answer (passi2019problem, ). Our results show that community members drew upon their lived experiences and questioned the motivation behind the application, such as challenging the police department’s claim of ‘efficient resource allocation,’ which they argued may unintentionally exacerbate racial profiling in low-income neighborhoods. Participants also question whether increased police presence truly leads to safer neighborhoods, further suggesting a need to reconsider the predominant focus on quality-of-life crimes.
@@
Here, community members also shared that there needed to be transparency regarding the use of such tools by government agencies but also that they needed to operate in the public interest; i.e. - improving the safety of citizens through fair policing practices that do not indiscriminately target and racially profile members of the community. Community members brought into question the street-level practices of law enforcement agents through which data is collected about low-income and minority communities. Similarly, technical participants, who are also community members, shared that such policing practices might perpetuate a vicious cycle by consistently collecting crime data in low-income neighborhoods. This cycle, or feedback loop, could result in increased policing and continuous data collection, a concern also raised in prior work (veale2018fairness, ). In line with recent work in the public sector (kuo2023understanding, ; stapleton2022has, ; zhang2023deliberating, ), our results provide further evidence that impacted community members are indeed indispensable to the AI deliberation process and can help avoid major pitfalls in AI systems. That is, involving impacted community members at the onset of the AI deliberation process can help identify critical fairness, accountability, transparency, and ethics (FATE) issues before such systems are developed and deployed. As highlighted by Kawakami et al. (kawakamistudying, ), this concern was also raised by public sector workers who believed that collaborating with community members could help address ”fundamental and root problems” that are missed otherwise.
@@
Here, the application acted as a boundary object (star1989structure, ) and helped facilitate detailed conversations with community members. The application helped scaffold nuanced and context-specific discussions with community members about policing practices in their city, and the use of AI systems by police departments, and helped articulate design guidelines for developing ethical AI systems that are centered in the needs of the impacted community.

### 19 Mar 2024 v2

@@ 4.2. Participants’ Interactions With Algorithmic Crime Mapping (RQ2)
We conducted in-lab activities with our participants where we first gave them an overview of the application and explained to them how they could change parameters to produce different crime maps. Next, we asked the participants to estimate the number of hotspots on the maps based on their knowledge of the city.
@@
We learned that LEAs used their hyperlocal knowledge of the different intersections (within neighborhoods) in the city to identify epicenters for criminal activity, however, in doing so, significantly overestimated the number of hotspots. For instance, an LEA who was interacting with the crime map shared that he was looking at all the major intersections in the vicinity of hotspots and tying criminal activity to his street-level knowledge of those intersections:
@@
LEAs could have been overestimating hotspots for several reasons. Evidence in criminology literature indicates that police officers may overestimate the prevalence of crime in certain neighborhoods or communities (Buil-Gil et al., 2021). Here, the perception of crime can be influenced by several factors such as the nature of policing work, media coverage of crime, and perceptions about low-income and minority communities (Smith and Alpert, 2007). There is also evidence that indicates that micro-level crime analysis as undertaken by LEAs (i.e., assessing specific intersections) is prone to a larger risk of selection and reporting bias as compared to crimes aggregated and assessed at a larger scale. However, in our conversations with the LEAs, we also learned that they did not draw clear distinctions between violent crimes and quality-of-life crimes. They often referred to the fact that different types of crimes can be interlinked and at times quality-of-life crimes can escalate to more serious crimes. That is, they drew on their domain expertise to focus on several hyperlocal hotspots instead of a general hotspot covering a neighborhood that a squad car could easily patrol. This finding was further supported by the next activity where we asked participants to place green circles on the map that demarcate the area that one patrol car should cover.
@@
During our observations of participants’ interactions with different crime maps, we recorded whether participants changed the parameters to create different maps and their responses to these different maps. Please see Appendix B for results of the quantitative analysis. Both technical participants and community members interacted with parameters and generated different maps. Some participants gamified the process for themselves and started to guess how changing certain parameters would affect the map. Technical participants shared that they had seen static geographical heatmaps before and knew how to generate them by writing code but had never interacted with an interactive map. Participant T9 shared:
@@
Several community members appreciated the opportunity to interact with such an application and understand how resources might be allocated using such tools. They shared that such transparency, even though carried out as a lab activity, helped them develop trust. For instance, a community member shared:
@@
Several community members also shared that they liked the easier maps because those maps provided them with bigger circles (i.e., hotspots) and gave them a quick overview of the neighborhoods. They shared that these would be a useful communication and awareness tool for the community.
The results, regarding LEAs, were quite surprising for the research team. We had anticipated that LEAs would be most likely to interact with the parameters, generate different maps, and connect the findings to street-level occurrences over different periods. However, none of the LEAs changed parameters and spent a significant amount of time interpreting the first map presented to them. Instead of using the crime maps to augment their knowledge of geospatial and temporal characteristics of different crimes, LEAs quickly started connecting their street-level domain knowledge to the hotspots on the map, i.e. - the first piece of information presented to them. Instead of using the tool as a decision aid that would augment their knowledge, they used it as a means to validate what they already knew about certain neighborhoods, sections of highways, and intersections. When urged to change parameters and create a different map, an LEA who had experience working with crime maps stated:
@@
The LEA makes an interesting point here regarding the data that we further delve into in the next subsection. However, for the most part, we witnessed that LEAs exhibited a strong anchoring bias where the default map (i.e., the ‘anchor’) provided a ‘satisfactory outcome’ that seemed to align with what they already knew about the city. Anchoring bias is a cognitive bias that refers to the tendency of practitioners to rely too heavily on the first piece of information encountered when making decisions and has been observed across several domains (Furnham and Boo, 2011; Azzopardi, 2021; Saposnik et al., 2016). However, as we discuss in the next section (see Section 4.3), the LEAs asked for features (e.g., the ability to draw districts on the map, see spatiotemporal relationships between different types of crime, and use information from on-the-ground investigations) that would have allowed them to derive more context-specific information.
@@
We wanted to assess participants’ level of engagement while interacting with the application and their mental workload as a result of this interaction. To accomplish this, we conducted the NASA-TLX survey and assessed participants on four factors (i.e., mental demand, temporal demand, performance, and frustration level). To assess the differences in scores across participants and their significance, we conducted quantitative analysis, the details of which are provided in Appendix C.
@@
The third parameter we considered was performance (PF) which determines the success of the users in carrying out their duties and how satisfied they were with the results of their work. Both the LEAs and technical participants scored high on this parameter which depicts that these participants were confident about their results. On the other hand, community members’ scores were in the medium range implying that they were uncertain about their results. As depicted in the previous section, both the LEAs and technical participants had some experience working with crime maps or geospatial heatmaps. However, technical participants interpreted maps based on their understanding of heatmaps whereas LEAs interpreted maps based on their street-level domain knowledge.
@@
Taken together, the results for LEAs are surprising. The NASA-TLX results show that LEAs were deeply invested in interpreting the maps presented to them (i.e., they scored high on mental and temporal demand), considered that they performed well when interpreting the hotspots and placing squad cars (i.e., scored high on performance) and did not find the task to be frustrating (i.e., scored low on frustration). However, as noted previously, none of the LEAs changed parameters to generate different maps. Here, LEAs were more interested in posthoc analysis and interpreting the information provided to them instead of generating new maps.
4.3. Participants’ Needs Regarding the Ethical Design & Use of AI Crime Mapping (RQ3)
@@
Several community members and technical participants found utility in algorithmic crime mapping as a public resource, however, questioned the underlying motivation of police departments themselves. Much of our conversations in this regard centered around ongoing policing practices where they questioned the trustworthiness of police reports by drawing on cases of police brutality where the police reports did not match the video evidence shared by bystanders on social media. For instance, participant C21 shared:
”From my personal experience, it seems that in black neighborhoods police are stopping more drivers per month per hour for all kinds of things, and for things that they don’t stop white drivers in the suburbs for. There are penalties for more and more people for using illegal drugs in the city, but people living in the suburbs use the same drug. But they get off with a warning and they think they are somehow different and too good for prosecution and they get some kind of a better deal. So there is a wide widespread perception in society, that law enforcement is not really fair. And on one hand, I would like to believe ’no, it really is fair’. But the evidence is coming up more and more in recent years. Like with the development of the cell phone, people take photographs and take videos and it doesn’t match the police reports. Before we had this technology, and some person who was arrested would say that the police beat him up and we would hear the police officer say that ‘he was resisting arrest’.” - C21, Community Member
This concern was shared by several community members who believed that such technologies could lead to racial profiling and amplify racial disparities. Participants shared examples of cases where the defendants were wrongfully convicted by association and forced into plea deals. Here, technical participants also brought into question the objectivity of the data analysis process itself and the need for public transparency and accountability mechanisms that keep such technologies in check. For instance, participant T5 shared the following when discussing the role of a data analyst working in a crime analysis lab:
@@
This concern was also echoed by other technical participants who shared examples of how data can be tweaked and cherry-picked and drew attention to the subjective nature of data science practice, an ongoing concern that has been raised in prior work (Mothilal et al., 2024; Das et al., 2024; Passi et al., 2017; Passi and Sengers, 2020). Finally, as discussed in Section 4.1, participants shared examples of differences in policing practices in urban and suburban neighborhoods, examples of speed traps, and people in low-income neighborhoods being set up for petty crimes to draw attention to policing practices that unethically target vulnerable communities instead of protecting and serving them. Here, community members sought accountability from police departments regarding street-level policing practices, whereas, technical participants wanted more transparency and accountability regarding data science practices that drive algorithmic crime mapping and data-driven policing.
@@ 4.3.2. Tensions Around the Ethical Basis of Data Collection
Several community members discussed the need for trustworthiness and accountability regarding data collection practices and discussed instances of falsehoods in police reports, differences in data that is available about urban versus suburban communities, and policing practices designed to target people (e.g., speed traps). Here, some participants also recognized that several government agencies collect data about citizens while conducting their daily operations but wanted to ensure that it is collected lawfully and with some oversight from the judicial system. For instance, a participant shared:
@@
This concern was also echoed by LEAs who shared that it was necessary to collect data legally and ethically otherwise it could not be used for investigations and would not be admissible in court during legal proceedings. An LEA shared:
@@
Here, the LEA provides an example of the police department getting access to people’s cell phone data in a specific neighborhood, however, states that they must make their case before a judge and get a court order approved before receiving access. LEAs also shared that their real-time crime mapping technology was based on service calls and not police reports and that people’s biases played a significant role regarding the calls they receive and investigate:
@@
Several LEAs recognized the concerns surrounding racial profiling but drew attention to the nature of their work where they were required to investigate service calls as well as patrol the zones that were assigned to them. Similarly, several LEAs shared the nuances of their street-level labor that we discuss next.
@@ 4.3.3. Nuances of Policing Practices
”I never knew the algorithm that was used - that was never shared. There were a couple of occasions where me and my team would be given a packet of information about a specific crime that was occurring in our area of responsibility, and sometimes the predictive information that was given to us wasn’t consistent with what we knew as law enforcement officers just being in these neighborhoods.” -L5, Law Enforcement Agent
@@
Here, it is interesting to note that both the technical participants and LEAs drew attention to the nature of data science practice. Technical participants discussed the elastic and subjective choices that data scientists make during analysis, whereas, LEAs discussed the need for data scientists to understand the nature of street-level discretionary work that they undertake. These themes have been observed across different contexts where domain experts have drawn attention to nuances of their professional practice that cannot be accounted for by quantitative data, and are generally unobservable to people outside the profession but significantly impact the decision-making process (Kawakami et al., 2022a; Clare Elish and Anne Watkins, 2020; Spektor et al., 2023).
@@ 5. Discussion
5.1. Eliciting Feedback from Community Members to Inform AI Problem Formulation
@@
Early deliberations regarding the development and use of AI systems often omit impacted stakeholders because of their lack of technical or data science knowledge which supposedly limits them from providing meaningful feedback that can inform systems design (Kuo et al., 2023; Delgado et al., 2023). However, as our results highlight, community members, irrespective of technical knowledge, are able to provide detailed feedback regarding whether the technical formulation of a problem is even socially relevant, draw attention to systemic and structural issues that might be invisible to AI developers, and help uncover the downstream impact of such systems on their communities. That is, impacted stakeholders have a critical role to play at the earliest stages of machine learning problem formulation where data scientists translate broad and nondescript objectives into tractable problems that machine learning can answer (Passi and Barocas, 2019). Our results show that community members drew upon their lived experiences and questioned the motivation behind the application, such as challenging the police department’s claim of ‘efficient resource allocation,’ which they argued may unintentionally exacerbate racial profiling in low-income neighborhoods. Participants also question whether increased police presence truly leads to safer neighborhoods, further suggesting a need to reconsider the predominant focus on quality-of-life crimes.
@@
Here, community members also shared that there needed to be transparency regarding the use of such tools by government agencies but also that they needed to operate in the public interest; i.e. - improving the safety of citizens through fair policing practices that do not indiscriminately target and racially profile members of the community. Community members brought into question the street-level practices of law enforcement agents through which data is collected about low-income and minority communities. Similarly, technical participants, who are also community members, shared that such policing practices might perpetuate a vicious cycle by consistently collecting crime data in low-income neighborhoods. This cycle, or feedback loop, could result in increased policing and continuous data collection, a concern also raised in prior work (Veale et al., 2018). In line with recent work in the public sector (Kuo et al., 2023; Stapleton et al., 2022b; Zhang et al., 2023), our results provide further evidence that impacted community members are indeed indispensable to the AI deliberation process and can help avoid major pitfalls in AI systems. That is, involving impacted community members at the onset of the AI deliberation process can help proactively identify critical fairness, accountability, transparency, and ethical (FATE) issues before AI systems are developed and deployed; a significant ongoing concern for Responsible AI (RAI) practices. As highlighted by Kawakami et al. (KAWAKAMI et al., 2023), this concern was also raised by public sector workers who believed that collaborating with community members could help address ”fundamental and root problems” that are missed otherwise.
@@
Here, the application acted as a boundary object (Star, 1989) and helped facilitate detailed conversations with community members. The application helped scaffold nuanced and context-specific discussions with community members about policing practices in their city, and the use of AI systems by police departments, and helped articulate design guidelines for developing ethical AI systems that are centered in the needs of the impacted community.